OpenStudy (anonymous):

If you have $1,000.00 on a credit card and the interest rate is 23.7% per year, how much interest will you pay for one year? (Multiply rather than compound your interest.) $2,370.00 $23.70 $237.00 $230.70

OpenStudy (stamp):

Identify. I = Prt P = $1,000 r = 23.7% t = 1 year Substitute. I = 1,000 * 23.7% * 1 Simplify. I = 1,000 * 0.237 * 1 I = 237 $237